OpenStudy (anonymous):

Solve x^2 + 9x + 9 = 0?

OpenStudy (anonymous):

You will need to use the quadratic formula for this one

OpenStudy (anonymous):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (vishal_kothari):

irrational roots...

OpenStudy (anonymous):

x^2+9x+81/4=-9-81/4 (x+9/2)^2=-117/4 x+9/2=+/_sqrt-117/2 the roots are imaginary

OpenStudy (anonymous):

x^2+9 x+9 = 0 Subtract 9 from both sides: x^2+9 x = -9 Add 81/4 to both sides: x^2+9 x+81/4 = 45/4 Factor the left hand side: (x+9/2)^2 = 45/4 Take the square root of both sides: sqrt(x+9/2)^2 = (3 sqrt(5))/2 Eliminate the absolute value: x+9/2 = -(3 sqrt(5))/2 or x+9/2 = (3 sqrt(5))/2 Subtract 9/2 from both sides: x = -3/2 (3+sqrt(5)) or x+9/2 = (3 sqrt(5))/2 Subtract 9/2 from both sides: x = -3/2 (3+sqrt(5)) or x = 3/2 (-3+sqrt(5))